Excess hydrochloric acid and nitrous acid are reacted with 45.0 g of aniline. The products are the explosive compound benzenediazonium chloride and water. The percentage yield for benzenediazonium chloride is 55.0%. Find the product obtained and the product expected.
1
Expert's answer
2020-04-27T01:52:18-0400

Solution:


The balanced chemical equation:

C6H5NH2 + NaNO2 + 2HCl = C6H5N2Cl + NaCl + 2H2O


Moles of aniline = Mass of aniline / Molar mass of aniline

Molar mass of aniline (C6H5NH2) is 93.13 g/mol.

Moles of aniline = n(C6H5NH2) = (45.0 g) / (93.13 g/mol) = 0.4832 mol.


According to the chemical equation:

Moles of aniline = Moles of benzenediazonium chloride.

n(C6H5NH2) = n(C6H5N2Cl) = 0.4832 mol.


Expected mass of benzenediazonium chloride = Moles of C6H5N2Cl × Molar mass of C6H5N2Cl

Molar mass of benzenediazonium chloride (C6H5N2Cl) is 140.57 g/mol

mexpected(C6H5N2Cl) = (0.4832 mol) × (140.57 g/mol) = 67.92 g

Obtained mass of C6H5N2Cl = % yield × Expected mass of C6H5N2Cl

mobtained(C6H5N2Cl) = (0.55) × (67.92 g) = 37.36 g


Answer:

The product expected is 67.92 g.

The product obtained is 37.36 g.